OpenStudy (anonymous):

How would I solve this problem? (2x^2+ 15x +8 )/(2x^2 + 7x +4)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

What do you want to solve?

OpenStudy (anonymous):

Factorise first

OpenStudy (anonymous):

Like I must simplify it .

OpenStudy (anonymous):

Then see what happens

OpenStudy (anonymous):

Yeah thats the part where I am stuck at. I am just not able to factor it out.

OpenStudy (anonymous):

Do you want metto do that 4 you

OpenStudy (anonymous):

Yes please but could you explain it too:).

OpenStudy (anonymous):

Pas de problem!

OpenStudy (anonymous):

use the quadratic formula to solve it...

OpenStudy (anonymous):

(2x^2+ 15x +8 )/(2x^2 + 7x +4) (2x^2+ 15x +8 ) to factorise .... multiply the 2 in 2x^2 by the + 8 and find a pair of factors that sum to -15 2 * -8 = -16 therefore factors are (1 and -16) now replace the -15 x with 1x (we just write x) and -6x 2x^2+ x -16x +8 ...now group and bracket and further factorise ... (2x^2+ x) -(16x +8) = x(2x+1)-8(2x+1) the brackets are identicval so their contents are a factor. The other factor is made up from the outside elements x and -8 so we have (2x^2+ 15x +8 ) = (2x+1)(x-8) Do the same with the other and you get ... (2x^2 + 7x +4) =(2x^2 + 8x -x+4) (2x^2 + 7x +4) =(2x+1)(x+4) now write the fraction in terms of factorised results (2x+1)(x-8) / (2x+1)(x+4) ...and you will see that (2x+1) factors out leaving (x-8) /(x+4)

OpenStudy (anonymous):

I am soooooo sorry I read -15 instead of +15 so my answer is incorrect.

OpenStudy (anonymous):

You will need to compleyte the square or use the quadratic formula as Shankvee suggests

OpenStudy (anonymous):

Ok thank you very much:)

OpenStudy (anonymous):

the equations clearly do not have integral roots as the discriminant is not a perfect square for both numerator and denominator so the only way is to use the quadratic formula x=((-b+sqrt(b^2-4ac))/2a where a,b,c are the coeffecients of x^2 , x and cis constant respectively...

OpenStudy (anonymous):

I really do appreciate it.

OpenStudy (anonymous):

Are you sure it is +15 and not -15 in the original question because that would make life very easy for you.

OpenStudy (anonymous):

Yes they are all plus signs. Still can you tell me how it would be if it was -15?

OpenStudy (anonymous):

no -15 makes no difference it should be a -8 or -2x^2

OpenStudy (anonymous):

if you have -8 thenyou can factorise numerator as 2x^2 +16x-x-8 2x(x+8)-1(x+8) (2x-1)(x+8)

OpenStudy (anonymous):

can you guys help me?

OpenStudy (anonymous):

yeah mosudan go ahead... You should ask a question really...

OpenStudy (anonymous):

7/12 + 2 2/3 -1.8

OpenStudy (anonymous):

7/12=0.58 2*2/3=1.33 so 0.58+1.33-1.8=0.08

OpenStudy (anonymous):

is there a way we can do this without a calculator ?